# ./references.bib

### v1

@@ 8 Applications to a line gapped system
Now we will provide an example of a physical system where the non-Hermiticity comes from a non-Hermitian Hamiltonian which is line gapped. Consider the generalized Haldane model over a bi-partite honeycomb lattice given in [haldane] with the following tight binding model hamiltonian;
@@
The first sum runs over all the lattice sites with $A$ and $B$ lattices having on-site energies ±$\pm M$. The second sum is kinetic energy with nearest neighbor coupling coefficient $t$. Lastly we have that the third sum is over next-nearest-neighbor pairs and has a direction-dependent phase factor $e^{i\phi(n_{\alpha},m_{\alpha})}$ a geometrically determined sign. When $\mu=0$ in the Hamiltonian, we have no lossy sites and the Hamiltonian is Hermitian. By making the model lossy, the Hamiltonian becomes non-Hermitian.
Our example tight binding model will be comprised of an inner topological insulator surrounded by a ring of trivial insulator which is then surrounded by a ring of a lossy trivial insulator, we call it the Haldane Heterostructure as in [cerjan_koekebier_Schulz_Baldes]. Again the loss comes from having on site energies being ±$\pm M-i\mu$ on the lossy Trivial insulator sites. The diagram for the tight binding model and unit cell is shown in Table 2. Furthermore, we will focus our attention on fixing the parameters for the tight binding model to the ones listed in Table 3.
@@
Trivial Insulator (Lossy) $0.5\sqrt{3}$ $0.2$ $1$ 00 N/A
Trivial Insulator (Non-Lossy) $0.5\sqrt{3}$ 00 $1$ 00 N/A
Topological Insulator 00 00 $1$ $0.5$ $\frac{\pi}{2}$
@@
The Hamiltonian will be non-Hermitian, and sparse for large system size. In some basis the diagonal entries will consist of complex (±$\pm M-i\mu$), purely real ±$\pm M$ and zeros as seen in Equation 16. In Equation 16, $C$ is a matrix with zeros on the diagonal.
@@
The 2d Clifford composite operator and the quadratic composite operator are as follows
@@
$Q^{\textup{2d}}_{(\kappa x,\kappa y,E)}(\kappa X,\kappa Y,H)$ $=\kappa\left(X-x\right)^{2}+\kappa\left(Y-y\right)^{2}+(H-E)^{\dagger}(H-E).$
@@
$\mu_{(\kappa\bm{x},0)}^{\text{Q}}\left(\kappa\bm{X},H\right)$
@@
$\left|\bar{\mu}_{(\kappa\bm{x},0)}^{\text{C}}\left(\kappa\bm{X},H\right)-\mu_{(\kappa\bm{x},0)}^{\text{Q}}\left(\kappa\bm{X},H\right)\right|$
@@
$\mu_{(\kappa\bm{x},1)}^{\text{Q}}\left(\kappa\bm{X},H\right)$
@@
$\left|\bar{\mu}_{(\kappa\bm{x},1)}^{\text{C}}\left(\kappa\bm{X},H\right)-\mu_{(\bm{x},1)}^{\text{Q}}\left(\kappa\bm{X},H\right)\right|$
@@
$\mu_{(\kappa\bm{x},i)}^{\text{Q}}\left(\kappa\bm{X},H\right)$
@@
$\left|\bar{\mu}_{(\kappa\bm{x},i)}^{\text{C}}\left(\kappa\bm{X},H\right)-\mu_{(\bm{x},i)}^{\text{Q}}\left(\kappa\bm{X},H\right)\right|$
Table 4: Shown are the Clifford linear gap (Figure 9,12, and 15), quadratic gap (Figure 10,13, and 16) as well as the difference between the gaps in Figure 11,14,and 17 of the Haldane Heterostructure with $(\bm{x},E)=(x,y,E)$ being the probe site with varying energy prove site $E=0,1,i$ along the rows of the table. Note that $(\bm{X},H)=(X,Y,H)$ are the two Position operators and Hamiltonian.
@@ 9 Conclusion
We have extended the theory of $\epsilon$-pseudospectrum [cerjan_loring_vides_2023] to allow for the study of systems that have more than one non-Hermitian operator. Along the way we have also proven some propositions about a special case of the non-Hermitian spectral localizer in [cerjan_koekebier_Schulz_Baldes] which we call the Clifford composite operator. We have also shown that even with perturbations to non-Hermitian $B_{i}$ the quadratic gap is local. With that said the benefits of the quadratic composite operator is that minimizing the corresponding gap function allows us to minimize eigen errors for various observables/matrices whithout increasing the computational complexity of the problem.

### v2

@@ 8 Applications to a line gapped system
Now we will provide an example of a physical system where the non-Hermiticity comes from a non-Hermitian Hamiltonian which is line gapped. Consider the generalized Haldane model over a bi-partite honeycomb lattice given in [13] with the following tight binding model hamiltonian;
@@
The first sum runs over all the lattice sites with $A$ and $B$ lattices having on-site energies $\pm M$. The second sum is kinetic energy with nearest neighbor coupling coefficient $t$. Lastly we have that the third sum is over next-nearest-neighbor pairs and has a direction-dependent phase factor $e^{i\phi(n_{\alpha},m_{\alpha})}$ a geometrically determined sign. When $\mu=0$ in the Hamiltonian, we have no lossy sites and the Hamiltonian is Hermitian. By making the model lossy, the Hamiltonian becomes non-Hermitian.
Our example tight binding model will be comprised of an inner topological insulator surrounded by a ring of trivial insulator which is then surrounded by a ring of a lossy trivial insulator, we call it the Haldane Heterostructure as in [3]. Again the loss comes from having on site energies being $\pm M-i\mu$ on the lossy Trivial insulator sites. The diagram for the tight binding model and unit cell is shown in Table 2. Furthermore, we will focus our attention on fixing the parameters for the tight binding model to the ones listed in Table 3.
@@
Trivial Insulator (Lossy) $0.5\sqrt{3}$ $0.2$ 11 0 N/A
Trivial Insulator (Non-Lossy) $0.5\sqrt{3}$ 0 11 0 N/A
Topological Insulator 0 0 11 $0.5$ $\frac{\pi}{2}$
@@
The Hamiltonian will be non-Hermitian, and sparse for large system size. In some basis the diagonal entries will consist of complex ($\pm M-i\mu$), purely real $\pm M$ and zeros as seen in Equation 18. In Equation 18, $C$ is a matrix with zeros on the diagonal.
@@
The 2d non-Hermitian spectral localizer and the quadratic composite operator are as follows
@@
$Q^{\textup{2d}}_{(\kappa x,\kappa y,E)}(\kappa X,\kappa Y,H)$ $=\left[\kappa\left(X-x\right)^{2}+\kappa\left(Y-y\right)^{2}\right]\otimes\begin{bmatrix}1&\\ &1\end{bmatrix}$
$\qquad+(H-E)^{\dagger}(H-E)\otimes\begin{bmatrix}1&\\ &0\end{bmatrix}+(H-E)(H-E)^{\dagger}\otimes\begin{bmatrix}0&\\ &-1\end{bmatrix}$.
@@
$\mu_{(\kappa\bm{x},0)}^{\text{RQ}}\left(\kappa\bm{X},H\right)$
@@
$\left|\bar{\mu}_{(\kappa\bm{x},0)}^{\text{C}}\left(\kappa\bm{X},H\right)-\mu_{(\kappa\bm{x},0)}^{\text{RQ}}\left(\kappa\bm{X},H\right)\right|$
@@
$\mu_{(\kappa\bm{x},1)}^{\text{RQ}}\left(\kappa\bm{X},H\right)$
@@
$\left|\bar{\mu}_{(\kappa\bm{x},1)}^{\text{C}}\left(\kappa\bm{X},H\right)-\mu_{(\bm{x},1)}^{\text{RQ}}\left(\kappa\bm{X},H\right)\right|$
@@
$\mu_{(\kappa\bm{x},i)}^{\text{RQ}}\left(\kappa\bm{X},H\right)$
@@
$\left|\bar{\mu}_{(\kappa\bm{x},i)}^{\text{C}}\left(\kappa\bm{X},H\right)-\mu_{(\bm{x},i)}^{\text{RQ}}\left(\kappa\bm{X},H\right)\right|$
Table 4: Shown are the Clifford linear gap (Figure 9,12, and 15), right quadratic gap (Figure 10,13, and 16) as well as the difference between the gaps in Figure 11,14,and 17 of the Haldane Heterostructure with $(\bm{x},E)=(x,y,E)$ being the probe site with varying energy prove site $E=0,1,i$ along the rows of the table. Note that $(\bm{X},H)=(X,Y,H)$ are the two Position operators and Hamiltonian.
@@ 9 Conclusion
We have extended the theory of $\epsilon$-pseudospectrum [4] to allow for the study of systems that have more than one non-Hermitian operator. Along the way we have also proven some propositions about a special case of the non-Hermitian spectral localizer in [3] which we call the non-Hermitian spectral localizer. We have also shown that even with perturbations to non-Hermitian $B_{i}$ the quadratic gap is local. With that said the benefits of the quadratic composite operator is that minimizing the corresponding gap function allows us to minimize eigen errors for various observables/matrices whithout increasing the computational complexity of the problem.